Laptops shangai computer purchased 20 laptops from a Chinese computer manufacturers shangai paid the bill after 45days paying a finance charges of 150 if the Chinese company charges 10%interest find the cost of 20 laptops
1
Expert's answer
2021-09-27T16:34:13-0400

Time period in years "=45days\/30days=1\/8"

Interest, I =PRT

"P=\frac{I}{R\times T}=\frac{150}{(10/100\times 1/8)}=150/0.0125=\$12,000"

The original cost of 20 laptops excluding interest is "\$12,000"